when asked her age, Miss Grundy refused to tell. After being begged for a hint, however, she finally admitted that in 12 years she hoped to br three times as old as she was 42 years ago.
how old is she now ?

Answer by htmentor(1343)

when asked her age, Miss Grundy refused to tell. After being begged for a hint, however, she finally admitted that in 12 years she hoped to br three times as old as she was 42 years ago.
==============
Let a = her age now
In 12 years her age will be a + 12
Her age 42 years ago = a - 42
And her age in 12 years is 3 times her age 42 years ago:
a + 12 = 3(a - 42)
Solve for a:
a + 12 = 3a - 126
2a = 138
a = 69
So she is 69 years old now.
Check:
69+12 = 81 = 3(69-42) = 3(27)